Calculate the percentage abundance of Br-80 and the average atomic weight of bromine in a brine sample that contains 50.54% Br-79 (78.918 amu) and Br-80 (80.916 amu). Express percent abundance up to 2 decimal places


1
Expert's answer
2022-02-18T08:43:04-0500

Solution:

Percentage abundance of Br-81 + Percentage abundance of Br-79 = 100%

Hence,

Percentage abundance of Br-81 = 100% - Percentage abundance of Br-79

Percentage abundance of Br-81 = 100% - 50.54% = 49.46%

Percentage abundance of Br-81 = 49.46%


Average atomic mass = f1M1 + f2M2 +… + fnMn, where fi is the fraction representing the natural abundance of the isotope and Mi is the mass number (weight) of the isotope.

Hence,

Average atomic mass of Br = f1M1 + f2M2 = (0.5054 × 78.918) + (0.4946 × 80.916) = 79.906 amu

Average atomic mass of Br = 79.91 amu


Answers:

The percentage abundance of Br-81 is 49.46%

The average atomic weight of bromine is 79.91 amu